An advertising firm claims that the average expenditure for advertising for their customers is at least $12,500 per year. You want to test this claim, so you randomly select 10 of their customers. The results are listed below. At a sig. level of .01 can you reject this claim? Explain.
13,445 11,220 10,157 13,217 9,990
12,125 10,116 14,350 13,590
12,100

Answer by ewatrrr(24785)

 
Hi
Ho: p >= 12.500 (claim)
Ha: p <12,500
-----------------------------------------
Sample:
mean = 12031
s = 1606.595
---
test stat:: z(12031) = (12031-12500)/[1606.6/sqrt(10)] = - 499/508.1 = -.9821
-----
p-value = P(z < -.9821) = .1630
Conclusiion:
Since the p-value is greater than .01, accept claim.
The test results support the claim.